OpenStudy (anonymous):

find the inverse function of: y= x/3 + 5

OpenStudy (slaaibak):

A way to do this: y = x/3 + 5 now swap the x and y, and make y the subject. x = y/3 + 5 so y = 3x - 15 is the inverse function